# Efficient Conditional Gradient Methods for Solving Stochastic Convex Bilevel Optimization Problems

### v1

@@ 2 Algorithms and convergence guarantees
$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t})^{\top}v\right\}$
@@
Result: sequences $\{x_{t},z_{t}\}_{t\geq 1}$.
Initialize $x_{0}\in X$;
for $t=0,1,2,\ldots$ do
       if $t=0\mod q$ then
            Compute
      else
             Compute
      Compute
$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t})^{\top}v\right\}$
$x_{t+1}$ $:=x_{t}+\alpha_{t}(v_{t}-x_{t})$
if $t\geq q$ then
            Compute
We emphasize that, unlike [4], both Algorithms 1 and 2 are anytime methods that do not require careful initialization step, preset stepsizes, or a priori choice of the number of total iterations $T$ based on target accuracies. This anytime property enables us to establish asymptotic convergence guarantees to the optimal solution, as formally stated in Theorems 3 and 4 in Appendix B. To establish convergence rates for these algorithms, we impose the following standard assumptions on the bilevel problem (1). In the following, all norms refer to the Euclidean norm.
@@
Our convergence rate guarantees are detailed in Theorems 2 and 1.
Theorem 1 (One-sample stochastic setting).
@@
Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ given by $\sigma_{t}:=\varsigma(t+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1/2)$. Under 1, for any $t\geq 1$, with probability at least $1-\delta$, we have
$F(z_{t})-F_{\operatorname{opt}}\leq O\left(t^{-(1/2-p)}\sqrt{\log\left(\frac{dt^{2}}{\delta}\right)}\right)\quad\&\quad G(z_{t})-G_{\operatorname{opt}}$ $\leq O\left(t^{-p}\sqrt{\log\left(\frac{d}{\delta}\right)}\right).$
Theorem 2 (Finite sum setting).
@@
Suppose the expectations defining $F$ and $G$ in Eq. 1 are from uniform distributions over finite sets of size $[n]$. Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$, and $\sigma_{t}:=\varsigma(\max\{t,q\}+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1)$ and $S=q$. Under 1, with probability at least $1-\delta$, for any $t\geq q$, it holds that
$F(z_{t})-F_{\operatorname{opt}}$ $\leq O\left(\max\left\{1,\frac{q\log(q)}{t}\right\}t^{-(1-p)}\sqrt{\log\left(\frac{t^{2}}{\delta}\right)}\right),$
$G(z_{t})-G_{\operatorname{opt}}$ $\leq O\left(t^{-p}\sqrt{\log\left(\frac{1}{\delta}\right)}\right).$
We present a proof sketch here; complete proofs are provided in Appendix A. Both results share a common structure, with key differences highlighted below.
Proof sketch of Theorems 1 and 2.
We employ the notation $\Phi_{t}(x):=\sigma_{t}F(x)+G(x)$. Let $\widehat{\nabla\Phi}_{t}=\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}$ be the gradient estimate at step $t$. Using standard analysis for conditional gradient methods with the stepsize $\alpha_{t}=2/(t+2)$, Lemmas 2 and 1 in Appendix A establish the inequality
$t(t+1)\left(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}})\right)$ $+\textstyle\sum_{i\in[t]\setminus[k]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F(x_{i})-F_{\operatorname{opt}})$
$\textstyle\leq(k+1)k\left(\Phi_{k}(x_{k})-\Phi_{k}(x_{\operatorname{opt}})\right)+2(\sigma_{0}L_{f}+L_{g})D^{2}(t-k)$ (3)
$\quad\textstyle+\sum_{i\in[t]\setminus[k]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$
for every $t\geq k$. Proposition 1 in Appendix A uses (3) to provide bounds for both the upper- and lower-level problems. Note that in the proof of Algorithm 1, we set $k=0$, while in in the proof of Algorithm 2, we set $k=q$. Nevertheless, the underlying idea remains essentially the same.
The key is to bound the last term $\sum_{i\in[t]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$, which is the error term from using noisy gradient estimates. To establish this bound with high probability, we leverage the concentration properties of the STORM estimator employed in Algorithm 1. Thanks to [4, Lemma 4.1], we obtain the following key probabilistic guarantee. For any iteration $i\in[t]$ and confidence level $\eta_{i}\in(0,1)$, with probability at least $1-\eta_{i}$, we (jointly) have
$\left\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\right\|$ $\leq O\left(\frac{\sqrt{\log(4d/\eta_{i})}}{(i+1)^{1/2}}\right)\quad\&\quad\left\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\right\|\leq O\left(\frac{\sqrt{\log(4d/\eta_{i})}}{(i+1)^{1/2}}\right),$
where the notation $O(\cdot)$ hides problem-specific constants $L_{f},L_{g},\sigma_{f},\sigma_{g},D$. Proposition 2 combines these inequalities for all $i\in[t]$ with $\eta_{i}=\delta/(i(i+1))$ and applies a union bound to conclude that with probability at least $1-\delta$, for any $t\geq 1$, we have
$\sum_{i\in[t]}i\left\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\right\|\leq O\left((t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}\right).$
Substituting this bound into (3), and using the convexity of $F$ as well as the feasibility of $x_{t}\in X$, implying the inequality $G(x_{t})-G_{\operatorname{opt}}\geq 0$, we obtain
$F(z_{t})-F_{\operatorname{opt}}\leq O\left(\frac{\sqrt{\log(8dt^{2}/\delta)}}{(t+1)^{1/2-p}}\right).$
This in turn yields the bounds on $G(z_{t})-G_{\operatorname{opt}}$. Details are presented in Lemmas 6, 4 and 5.
Algorithm 2 uses the SPIDER estimator, resulting in slightly different probabilistic bounds that depend on the frequency $q$ in which we compute the full gradient and the mini-batch size $S$. Thanks to [41, Theorem 3.5], we show in Lemma 7 that with probability at least $1-\eta_{i}$, we (jointly) have
$\left\|\widehat{\nabla F}_{i}\!-\!\nabla F(x_{i})\right\|\leq O\left(\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)~{}~{}\&~{}~{}\left\|\widehat{\nabla G}_{i}\!-\!\nabla G(x_{i})\right\|\leq O\left(\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)$
when $1\leq i\leq q$, and
$\left\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\right\|$ $\leq O\left(\sqrt{\frac{(i-s_{i})}{S(i+1)(s_{i}+1)}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)$
$\left\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\right\|$ $\leq O\left(\sqrt{\frac{(i-s_{i})}{S(i+1)(s_{i}+1)}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right),$
when $i>q$, where $s_{i}:=q\lfloor i/q\rfloor$ is the closest multiple of $q$ below $i$. Since the numerator of the fraction term $i-s_{i}$ is always bounded by $q$, the same aggregation trick as in Proposition 3, except ignoring the first $q$ terms (which are handled in Lemma 8), leads to an improved dependence. Specifically, we can show that, with probability $1-\delta$, we have
$\sum_{i\in[t]\setminus[q]}i\left\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\right\|\leq O\left((8\sqrt{2}t-4q)\sqrt{\log\left({16t^{2}}/{\delta}\right)}\right).$
This improvement carries through to get the improved dependence on $t$ in Theorem 2; full details are given in Lemmas 8, 9, 10 and 11. ∎
Remark 1 (Sample complexity).
To attain an $\epsilon$-level optimality gap for both the outer- and inner-level problems, Algorithm 1 with $p=1/4$ requires $O(1/\epsilon^{4})$ stochastic oracle calls under the assumptions of Theorem 1. Likewise, under the assumptions of Theorem 2 and with $S=q=\lfloor\sqrt{n}\rfloor$ and $p=1/2$, Algorithm 2 achieves the same accuracy with $O(\sqrt{n}/\epsilon^{2})$ stochastic oracle calls.
@@ 3 Numerical results
We showcase the performance of our proposed algorithms in an over-parametrized regression and an $\ell_{1}$-constrained logistic regression problems. Each experiment is repeated $10$ times, and we report the average performance within a $4$-minute execution time limit. For IR-SCG, presented in Algorithm 1, we set $\sigma_{t}=\varsigma(t+1)^{-1/4}$, where $\varsigma=10$ in the first experiment and $\varsigma=0.01$ in the second experiment, along with $\alpha_{t}=2/(t+2)$. For IR-FSCG, presented in Algorithm 2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q\}+1)^{-1/2}$, where $\varsigma=10$ in the first experiment and $\varsigma=0.01$ in the second experiment, along with $\alpha_{t}=2/(t+1)$ for every $t\geq q$ and $\alpha_{t}=\log(q)/q$ for every $t<q$. In both experiments, the linear optimization oracle admits an analytical solution; see [25, Section 4.1].
For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1] with stepsize $\gamma_{t}=0.01/(t+1)$, SBCGF [4, Algorithm 2] with constant stepsize $\gamma_{t}=10^{-5}$ (both using $K_{t}=10^{-4}/\sqrt{t+1},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG [26] with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-7}/(t+1)^{3/4}$, $\rho_{t}=10^{3}(t+1)^{1/4}$, $r=0.5$), and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [20] with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=10^{-6}$) stepsizes. We initialized all algorithms with randomized starting points. For SBCGI and SBCGF, we generated the required initial point $x^{\prime}_{0}$ by running the SPIDER-FW algorithm [57, Algorithm 2] on the inner-level problem in (6), using a stepsize of $\gamma_{t}=0.1/(t+1)$. The initialization process terminated either after $10^{5}$ stochastic oracle queries or when the computation time exceeded 100 seconds, with the resulting point serving as $x_{0}$ for both SBCGI and SBCGF. The implementation of these algorithms rely on certain linear optimization or projection oracles. For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{1}$-norm ball intersecting with a half-space is required, which we employ CVXPY [10] to solve the problems, similar to the implementation in [4, Appendix F.1]. To compute the projection onto the feasible set required for aR-IP-SeG, we used [12, Algorithm 1]. To approximate the outer- and inner-level optimal values $F_{\operatorname{opt}},G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem and the reformulation (2) of the bilevel problem. All implementation details follow those in the corresponding papers.
@@ 3.1 Over-parameterized regression
We study the following bilevel optimization problem
$\begin{array}[]{cl}\displaystyle\min_{x\in{\mathbb{R}}^{d}}&\displaystyle F(x):=\frac{1}{|\mathcal{D}_{\text{val}}|}\|A_{\text{val}}x-b_{\text{val}}\|_{2}^{2}\\ \text{s.t.}&\displaystyle x\in\operatorname*{arg\,min}_{\|z\|_{1}\leq\delta}G(z):=\frac{1}{|\mathcal{D}_{\text{tr}}|}\|A_{\text{tr}}z-b_{\text{tr}}\|_{2}^{2},\end{array}$ (6)
@@
Figure 1 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) optimality gaps in terms of execution time. Our proposed stochastic methods outperform existing approaches. In particular, IR-FSCG achieves high-accuracy solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGI. Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI. We also observe that both SBCGI and SBCGF exhibit degradation in their inner-level optimality gaps over time. This behavior aligns with the theoretical guarantees in [4], where the inner-level convergence can only be bounded by $\epsilon_{g}/2$ without monotonic improvement guarantees. Furthermore, neither aR-IP-SeG nor SDBGD makes significant progress in optimizing the outer-level objective function. Finally, the table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm. Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free approach in [4]. We also observe that SDBGD and IR-SCG perform a similar number of iterations, as the $\ell_{1}$-norm projection oracle has comparable computational cost to the linear minimization oracle.
@@
Figure 1: The inner-level optimality gap (left) and the outer-level absolute optimality gap (right) over time for different algorithms on the over-parametrized regression problem Eq. 6.
3.2 Minimal norm logistic regression
Pedregosa et al. [39] studied the following $\ell_{1}$-constrained logistic regression problem
$\operatorname*{arg\,min}_{\|x\|_{1}\leq\beta}\frac{1}{n}\sum_{i\in[n]}\psi(a_{i}^{\top}x,b_{i})+\frac{\lambda}{2}\|x\|_{2}^{2},$ (7)
where $\psi$ is the logistic loss function. In this section we consider the bilevel variant of problem (7), in which we seek the minimal $\ell_{2}$-norm solution to $\ell_{1}$-constrained logistic regression problem, that is, we aim to solve the following bilevel optimization problem
$\min_{x\in{\mathbf{R}}^{d}}$ $\quad F(x):=\frac{1}{2}\|x\|_{2}^{2}$ (8)
s.t. $\quad x\in\operatorname*{arg\,min}_{\|z\|_{1}\leq\beta}G(z):=\frac{1}{n}\sum_{i\in[n]}\psi(a_{i}^{\top}z,b_{i}).$
We note that here we treat the outer-level as deterministic. Therefore, the randomness in the implemented algorithms is inherited from the solely the inner-level problem. We evaluated the method on the Madelon dataset [21], which consists of a data matrix $A\in\mathbb{R}^{n\times d}$ with $n=2000$ samples and $d=500$ features, along with an output vector $b\in\mathbb{R}^{n}$. We set $\beta=20$ for all experiments.
Figure 2 illustrates the convergence behavior of various algorithms in terms of execution time, showing the optimality gaps for the inner-level (left) and outer-level (right) problems. The table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm. The overall performance trends are consistent with the previous experiment. However, all baseline stochastic methods fail to make meaningful progress within the fixed time limit. In contrast, our proposed algorithms, in particular IR-SCG, demonstrate strong performance on both the lower- and upper-level problems.

### v2

@@ 2 Algorithm and convergence results for the one-sample stochastic setting
$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{\left(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}\right)^{\top}v\right\}$
@@
To establish convergence rates for Algorithm˜1, we impose the following standard assumptions on the simple bilevel problem (1). In the following, all norms refer to the Euclidean norm.
@@
First, we present the result for the case that $F$ is convex.
Theorem 1.
@@
Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and regularization parameters $\sigma_{t}:=\varsigma(t+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1/2)$. Under ˜1, for any $t\geq 1$, with probability at least $1-\delta$, it (jointly) holds that
$F(z_{t})-F_{\operatorname{opt}}\leq O\left(t^{-(1/2-p)}\sqrt{\log\left({dt^{2}}/{\delta}\right)}\right),\quad G(z_{t})-G_{\operatorname{opt}}$ $\leq O\left(t^{-p}\sqrt{\log\left({d}/{\delta}\right)}\right).$
Moreover, with probability $1$, we have
Theorem˜1 provides the formal convergence guarantees for the IR-SCG algorithm in the convex setting. If we set $p=1/4$, Algorithm˜1 achieves an $\epsilon$-level optimality gap for both the outer and inner problems with a sample complexity of $O(1/\epsilon^{4})$.
We next extend our analysis to the more general case where $F$ is not necessarily convex. In this setting, we require a different measure for stationarity. We define our stationarity measure using the Frank-Wolfe gap, which for a convex function $\mathcal{F}(x)=0$ if and only if $x$ is optimal. This provides a natural generalization for the non-convex case. Accordingly, for all $x\in X$, we define the functions
$\mathcal{F}(x):=\max_{v\in X_{\operatorname{opt}}}\nabla F(x)^{\top}(x-v),\quad\mathcal{G}(x):=G(x)-G_{\operatorname{opt}}.$ (3)
To avoid clutter, we present the result using the optimally tuned parameters of Algorithm˜1.
Theorem 2.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with stepsizes $\alpha_{t}=(t\!+\!1)^{-\omega}$ and regularization parameters $\sigma_{t}=\varsigma(t\!+\!1)^{-p}$ for any $t\!\geq\!0$, with $p\!=\!\frac{2}{7}$ and $\omega\!=\!\frac{6}{7}$. If Assumption 1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq 0$, it (jointly) holds that
$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}\leq O\left(t^{-\frac{1}{7}}\log\left({dt^{2}}/{\delta}\right)\right),\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{7}}\log\left({dt^{2}}/{\delta}\right)\right).$
Moreover, with probability $1$, we have
$\liminf_{t\to\infty}\mathcal{F}(x_{t})=0,\quad\lim_{t\to\infty}{\cal G}(x_{t})=0.$
3 Algorithm and convergence results for the finite-sum setting
In this section, we address the case where the expectations in (1) are finite sums over $n$ components. We introduce the Iteratively Regularized Finite-Sum Conditional Gradient (IR-FSCG) method, summarized in Algorithm˜2, where $n]:=\{1,\dots,n\}$. This approach is a conditional gradient method that leverages the SPIDER estimator, achieving variance reduction through periodic gradient recomputations with mini-batches of size $q$. While the SPIDER estimator has been previously integrated into projection-free methods [48, 57, 59, 22], here we show how it can be effectively adapted to our regularized setting. The primary advantage is that, unlike the method of Cao et al. [4], IR-FSCG preserves the desirable anytime properties of our approach while achieving a faster convergence rate that capitalizes on the finite-sum structure. Additionally, it does not require careful initialization, preset step sizes, or knowledge of the total number of iterations $T$ in advance.
To establish its convergence, we again rely on ˜1, noting that condition ˜1(e) holds automatically in this finite-sum setting. First, we present the result when $F$ is convex.
Result: sequences $\{x_{t},z_{t}\}_{t\geq 1}$.
Initialize $x_{0}\in X$;
for $t=0,1,2,\ldots$ do
 if $t=0\mod q$ then
     Compute
 else
      Compute
  Compute
$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{\left(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}\right)^{\top}v\right\}$
$x_{t+1}$ $:=x_{t}+\alpha_{t}(v_{t}-x_{t})$
if $t\geq q$ then
     Compute
Theorem 3.
@@ 3 Algorithm and convergence results for the finite-sum setting
Suppose $F$ and $G$ in Eq.˜1 are expectations over uniform distributions on $[n]$. Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$, and $\sigma_{t}:=\varsigma(\max\{t,q\}+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1)$ and $S=q$. Under ˜1, with probability at least $1-\delta$, for any $t\geq q$, it holds that
$F(z_{t})-F_{\operatorname{opt}}\leq O\left(c(t,q)\,t^{-(1-p)}\sqrt{\log\left({t^{2}}/{\delta}\right)}\right),\quad G(z_{t})-G_{\operatorname{opt}}\leq O\left(t^{-p}\sqrt{\log\left({1}/{\delta}\right)}\right),$
where $c(t,q)=\max\left\{1,{q\log(q)}/{t}\right\}$. Moreover, with probability $1$, we have
Theorem˜3 provides the formal convergence guarantees for the IR-FSCG algorithm in the convex setting. Moreover, when we set $S=q=\lfloor\sqrt{n}\rfloor$ and $p=1/2$, Algorithm˜2 achieves an $\epsilon$-level optimality gap for both the outer and inner problems with a sample complexity of $O(\sqrt{n}/\epsilon^{2})$. Besides, Algorithm˜2 can be readily extended when the outer- and inner-level problems involve different numbers of component functions ($n_{f}\neq n_{g}$) by introducing $(S_{f},q_{f})$ for the outer level and $(S_{g},q_{g})$ for the inner level. Details are omitted for brevity.
We next extend our analysis to the more general case where $F$ is not necessarily convex, using the auxiliary function (3) to assess stationarity. To avoid clutter, we present the result using the optimally tuned parameters of Algorithm˜2.
Theorem 4.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with stepsizes $\alpha_{t}=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$, and regularization parameters $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-p}$ with $p=\frac{1}{2}$ and $w=\frac{3}{4}$. If ˜1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq q$, it (jointly) holds that
$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}\leq O\left(t^{-\frac{1}{4}}\log\left({dt^{2}}/{\delta}\right)\right),\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{4}}\log\left({dt^{2}}/{\delta}\right)\right),$
where
$\beta_{i}=\begin{cases}\displaystyle\frac{\sigma_{0}(q+1)^{\omega}\log(q+1)}{q+1}\left(1-\frac{\log(q+1)}{q+1}\right)^{q-i}&0\leq i\leq q,\\ \sigma_{i}&i>q.\end{cases}$ (4)
Moreover, with probability $1$, we have
$\liminf_{t\to\infty}\mathcal{F}(x_{t})=0,\quad\lim_{t\to\infty}{\cal G}(x_{t})=0.$
@@ 4 Numerical results
We showcase the performance of our proposed algorithms in an over-parametrized regression and a dictionary learning problems. For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1], SBCGF [4, Algorithm 2], aR-IP-SeG [27], and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [21]. All implementation details follow those in the corresponding papers and available in Appendix˜D. To ensure reproducibility, all source codes are made available at https://github.com/brucegiang/CG-StoBilvl.
@@ 4.1 Over-parameterized regression
We first consider a simple bilevel optimization problem with a convex outer-level objective function
$\begin{array}[]{cl}\displaystyle\min_{x\in{\mathbf{R}}^{d}}\displaystyle F(x):=\frac{1}{|\mathcal{D}_{\text{val}}|}\|A_{\text{val}}x-b_{\text{val}}\|_{2}^{2}\quad\text{s.t.}\quad\displaystyle x\in\operatorname*{arg\,min}_{\|z\|_{1}\leq\delta}G(z):=\frac{1}{|\mathcal{D}_{\text{tr}}|}\|A_{\text{tr}}z-b_{\text{tr}}\|_{2}^{2},\end{array}$ (6)
@@
We evaluate the performance of different algorithms across $10$ experiments with random initializations and report their average performance within a $4$-minute execution limit. Figure˜1 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) optimality gaps in terms of execution time. Our proposed stochastic methods outperform existing approaches. In particular, IR-FSCG achieves high-accuracy solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGF. Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI. We also observe that both SBCGI and SBCGF exhibit degradation in their inner-level optimality gaps over time. This behavior aligns with the theoretical guarantees in [4], where the inner-level convergence can only be bounded by $\epsilon_{g}/2$ without monotonic improvement guarantees. Furthermore, neither aR-IP-SeG nor SDBGD makes significant progress in optimizing the outer-level objective function.
The table below also reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm. Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free approach in [4]. We also observe that SDBGD and IR-SCG perform a similar number of iterations, as the $\ell_{1}$-norm projection oracle has comparable computational cost to the linear minimization oracle.
@@
Figure 1: The inner-level optimality gap (left) and the outer-level absolute optimality gap (right) over time for different algorithms on the over-parametrized regression problem Eq.˜6.
4.2 Dictionary learning
$\begin{aligned} \min_{D,X}&\quad\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Dx_{i}\|_{2}^{2}\\ \text{s.t.}~&\quad D\in{\mathbf{R}}^{m\times p},\ X=[x_{1}\cdots x_{n}]\in{\mathbf{R}}^{p\times n}\\ &\quad\|d_{j}\|_{2}\leq 1,\quad\forall j\in[p],\quad\|x_{i}\|_{1}\leq\delta,\quad\forall i\in[n],\end{aligned}$
where we refer to $X$ as the coefficient matrix. In practice, data points usually arrive sequentially and the representation evolves gradually. Hence, the dictionary must be updated sequentially as well. Assume that we already have learned a dictionary $\hat{D}\in{\mathbf{R}}^{m\times p}$ and the corresponding coefficient matrix $\hat{X}\in{\mathbf{R}}^{p\times n}$ for some data set $A$. As a new dataset $A^{\prime}=\{a_{1}^{\prime},\dots,a_{n^{\prime}}^{\prime}\}\subseteq{\mathbf{R}}^{m}$ arrives, we intend to enrich our dictionary by learning $\tilde{D}\in{\mathbf{R}}^{m\times q}$ ($q>p$) and the coefficient matrix $\tilde{X}\in{\mathbf{R}}^{q\times n^{\prime}}$ for the new dataset while maintaining good performance of $\tilde{D}$ on the old dataset $A$ as well as the learned coefficient matrix $\hat{X}$.
This leads to the following simple bilevel problem with a nonconvex outer-level objective function
$\min_{\tilde{D},\tilde{X}}$ $\quad F(\tilde{D},\tilde{X})=\frac{1}{2n^{\prime}}\sum_{i\in[n^{\prime}]}\|a_{i}^{\prime}-\tilde{D}\tilde{x}_{i}\|_{2}^{2}$ (7)
s.t. $\quad\tilde{D}\in{\mathbf{R}}^{m\times q},\ \tilde{X}=[\tilde{x}_{1}\cdots\tilde{x}_{n}]\in{\mathbf{R}}^{q\times n^{\prime}},\quad\|\tilde{x}_{i}\|_{1}\leq\delta,\quad\forall i\in[n^{\prime}],$
$\quad\tilde{D}\in\operatorname*{arg\,min}_{\|z_{j}\|_{2}\leq 1,\forall j\in[q]}g(Z)=\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Z\hat{x}_{i}\|_{2}^{2},$
where we denote $\hat{x}_{i}$ as the prolonged vector in ${\mathbf{R}}^{q}$ by appending zeros at the end. We consider problem (7) on a synthetic dataset with a similar setup to [29, 4].
We evaluate the performance of different algorithms across $10$ experiments with random initializations and report their average performance within a $1$-minute execution limit. Figure˜2 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) problems in terms of execution time. Our proposed stochastic methods outperform existing approaches. In particular, IR-FSCG achieves better solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGI. Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI. We again observe that both SBCGI and SBCGF exhibit degradation in their inner-level optimality gaps over time.
The table below reports the average number of iterations and stochastic oracle calls over the $1$-minute time limit for each algorithm. Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free methods in [4]. We also observe that SDBGD performs fewer iterations than IR-SCG, as it relies on a more costly projection oracle.